OpenStudy (anonymous):

Find the expression if g(x)=x^2+2x: g(3)+g(7)

OpenStudy (waleed_imtiaz):

g(3)=(3)^2+2(3)=9+6=15 g(7)=(7)^2+2(7)=49+14=63 so 63+15=78

OpenStudy (anonymous):

\[[3^{2}+2(3)]+[7^{2}+2(7)]\] \[[9+6]+[49+14]\] \[15+63\] \[78\]